OpenStudy (anonymous):

x^2-14x+49 which is the correst answer?

OpenStudy (anonymous):

a=(x-7)^2 b=(x+7)^2 c=(x+14)^3 d=x+14

OpenStudy (haseeb96):

(a+b)^2=a^2-2ab+b^2 So the correct answer is A. ( (x-7)^2=x^2-14x+49